A buffer solution is made from 0.650 mol/L ammonia and 0.350 mol/L ammonium chloride (to total 1L of liquid). What is the pH of this buffer?

(Kb = 1.8 x 10−5)
1
Expert's answer
2020-09-14T08:17:54-0400

When a weak base is mixed with its salt, an alkaline buffer mixture is obtained.

For an alkaline buffer mixture, pH = 14-PKB + lg (Pine/SSOL)

RKB of ammonia = - lg Kb

Kb (NH4OH)=1,8∙10^-5

pH =14 + lg(1.8×10^-5) + lg (0.65 / 0.35) =9.5241=9.52